OpenStudy (anonymous):

Find the average cost of 50 units of a commodity whose cost function is given by C(x)=2.6x²-1.3x+5

OpenStudy (anonymous):

what is x??

OpenStudy (anonymous):

x is an alphabet of english language:)

OpenStudy (anonymous):

lol....

OpenStudy (anonymous):

if x is unit here then put c(50)=2.6*50^2-1.3*50+5 =6 440 average=6440/50=128.8

OpenStudy (anonymous):

hello tina_ann...........